A company has a capital structure that consists of $20 million of debt and $20 million of common equity, based upon current market values. The company’s yield to maturity on its bonds is 8%, and the current stock price is $35, the last dividend paid was $1.10 and the dividends are expected to grow at constant rate of 5% for long time. If the tax rate is 40%, what is this company's WACC assuming that there won’t be any new equity issuance?

6.55%

10.73%

8.51%

7.13%

9.34%
Answer #1

After-tax cost of debt=8*(1-tax rate)

=8*(1-0.4)=4.8%

Cost of equity=(D1/Current price)+Growth rate

=(1.1*1.05)/35+0.05

=8.3%

Total value=(20+20)=$40 million

WACC=Respective costs*Respective weight

=(20/40*4.8)+(20/40*8.3)

=6.55%